A 50.00 mL sample of 0.0950 M acetic acid (Ka = 1.8 x 10-5) is being titrated with 0.106 M NaOH.

What is the pH at the half-way point of the titration? (22.41 mL of 0.106 M NaOH has been added)

A) 3.06

B) 5.04

C) 4.44

D) 3.18

E) 4.74
Answer #1

In the halfway titration, we always have a buffer

pH = pKa + log(conjguate/acid)

and since it is half way then

conjugate = acid

so

pH = pKa + log(1) = pKa

pH = pKa

pKa = -log(Ka) = -log(1.8*10^-5) = 4.75

neareswt answer is E pH = 4.74